OpenStudy (anonymous):

Solve the initial value problem\[u''+6u'+12u=0\]with \(u(0)=1\) and \(u'(0)=0\)

OpenStudy (anonymous):

use characteristic equation a^2 + 6a+12 solve for a

OpenStudy (abb0t):

this is a second order homogeneous differential equation. do you know the general solution? \[y_p = e^{rt}c_1+e^{rt}c_2\]

OpenStudy (abb0t):

@modphysnoob is correct. except he used "a" so your solution would be: \[y_p = e^{at}c_1+e^{-at}c_2\]

OpenStudy (abb0t):

solve the quadratic

OpenStudy (anonymous):

I actually have the answer but I am just a little confused about how to do it... It's a spring mass problem and the answer is \[u=e^{-3t}cos\sqrt{3}t+\sqrt{3}e^{-3t}sin\sqrt{3}t\]

OpenStudy (abb0t):

then, plug in the initial conditions. take a few derivatives and violê

OpenStudy (abb0t):

solve for u(0)=1 take the derivative of your particular solution and plug in u(0)=0

OpenStudy (anonymous):

ok. when I factor \(a^2+6a+12=0\) I get \((-i a+\sqrt{3}-3 i) (i a+\sqrt{3}+3 i)\)

OpenStudy (anonymous):

that mean it is sinosoid

OpenStudy (anonymous):

which means it's \((\sqrt{3}\pm(3+a)i)=0\) how do i make it look like the general solution?

OpenStudy (anonymous):

can you help? because I'm so confused

OpenStudy (anonymous):

hold on, my computer is apparently running on steam engine

OpenStudy (anonymous):

a=-3 (+-) i* sqrt(3) which is \[Ae^{(-3+ i \sqrt{3})t}+Be^{(-3- i \sqrt{3})t}\]

OpenStudy (anonymous):

use euler identity and you will get A sin( ) + B cos ()

OpenStudy (anonymous):

now i can see it's \[Ae^{-3}sin(\sqrt{3}t)+Be^{-3}cos(\sqrt{3}t)\]then all i have to do is plug in the initial values, right?

OpenStudy (anonymous):

you are missing t e^(-3t)

OpenStudy (anonymous):

oh yeah. ok there's a t.

OpenStudy (anonymous):

\[u=e^{−3t}cos\sqrt{3}t+\sqrt{3}e^{−3t}sin\sqrt{3}t\]one more question. how do i determine if this equation is over damped, underdamped, or critically damped?

OpenStudy (anonymous):

so sin() and cos(x) just go up and down at t go on

OpenStudy (anonymous):